OpenStudy (anonymous):

If A is idempotent, show that I-2A is equal to I-2A inverse.

OpenStudy (anonymous):

Consider \[(I-2A)^{2} = I-4A+4A^{2}\] Because A is idempotent, \[A=A^{2}\] so our equation becomes\[I-4A+4A = I\] If something squared is the identity, it is its own inverse.